# Meta-reinforcement learning with minimum attention

### v2

@@ 4.2 The details of the algorithm implementation
Figure 2: (a) Training of Half-Cheetah. There is a reduction of variance from minimum attention, (b) Comparison of learning curves among model-free RL (SAC, (haarnoja2018)) and model-based RL (MB-MPO, (clavera2018), the proposed MB-MPO + minimum attention, weighting factor $\alpha=0.05$ and 11). The experiment is run with 10 episodes.
@@
Half-Cheetah Hopper Walker2D Humanoid
MB-MPO training 6692 $\pm$ 318 2475 $\pm$ 3.81 2399 $\pm$ 694 575 $\pm$ 177
MB-MPO meta-testing 6356 $\pm$ 132 467 $\pm$ 100 523 $\pm$ 174 315 $\pm$ 67
Table 1: Comparison of the model-based RL (MB-MPO) with regularization of minimum attention in meta-training and testing. Total reward at the epoch of 200K. For Hopper, the thigh is paralyzed. For Walker2D, the right thigh is paralyzed.
4.3 Dependency on the ensemble model size and imaginary sample size
Figure 4 shows the influence of the ensemble model size of $M$ and the imaginary sample size of $N$. When the ensemble model size of $M$ increases to 20, the total reward also increases. The imaginary sample size $N$ provides the benefit of increased total reward to $N=128$, but decreases when $N$ is increased to 256.
Figure 4: Half-Cheetah: Dependency on the ensemble model size and imaginary rollouts. When the ensemble model size of $M$ increases to 20, the total reward also increases. The imaginary sample size $N$ provides the benefit of increased total reward to $N=128$, but decreases when $N$ is turned to 256.
4.4 Comparison with state-of-the-art RL algorithms
The comparision is mostly based on the performance on standard locomotion tasks. We consider a model-free RL, SAC (Soft Actor-Critic) (haarnoja2018) and a model-based RL, MB-MPO (Model-Based Meta-Policy Optimization) (clavera2018). Firstly, we show the training of Half-Cheetah (Figure 2) in comparison of learning curves among the model-free RL (SAC), the model-based RL (MB-MPO), and the proposed MB-MPO with minimum attention, where the weighting factor is $\alpha=0.01$, $0.05$, and $1$. We have experienced a statistically significant reduction of variance from minimum attention. To demonstrate the generality of our approach, Figure 3 presents the learning curves for Hopper, Walker2D, and Humanoid. Across all three kinematically distinct agents, our method (MB-MPO + minimum attention, blue) consistently converges to a higher average return than the vanilla MB-MPO baseline (red). For the complex Humanoid task, our method also exhibits noticeably lower variance, indicating a more stable and reliable learning process.
The outperformance of MB-MPO with minimum attention is quantified in numbers in Table 1 and 2. Table 2 provides a detailed quantitative summary of the final performance on HalfCheetah after 200K iterations. In the standard meta-training setting, the best minimum attention factor ($\alpha=1.0$) achieves a 45% higher reward than MB-MPO, while simultaneously reducing the average feedback norm (Jacobian) by 28% and energy consumption by 15%. This demonstrates that the regularization does not trade performance for smoothness, but rather enables the discovery of policies that are both higher-performing and more efficient. Crucially, in the meta-testing scenario where the agent must adapt to a crippled leg, the performance benefits of our method are maintained or even amplified. The half-cheetah with minimum attention achieves a higher reward ($6822\pm 109$) than the vanilla baseline ($6356\pm 132$) and does so with a lower feedback norm and feedforward norm (jerk). This suggests that the smoother policies learned through minimum attention provide a better prior for rapid adaptation to significant out-of-distribution embodiment changes.
@@ 4.4 Comparison with state-of-the-art RL algorithms
We first establish the performance of our method on standard locomotion tasks. Figure 2 presents the learning curves on HalfCheetah, comparing our method against model-based (MB-MPO) and model-free (SAC) baselines. Our approach not only matches the sample efficiency of MB-MPO but also converges to a significantly higher average reward. Notably, the learning curves of our method also exhibit reduced variance across episodes, suggesting a more stable and reliable learning process, a key benefit for safe RL. This trend of superior performance and stability generalizes across other agents, as shown by the learning curves for Hopper, Walker2D, and Humanoid in Figure 3. Quantitative summary for all tested environments can be found in Appendix B.
@@ 4.5 Analysis of learned control strategies of minimum attention
Figure 6: Heatmap of feedback. With Half-Cheetah, the feedback ($\partial u/\partial x||^{2}$) in the projection of the states to a) X velocity and Z coordinates, b) Z velocity and X velocity, and c) Y angular velocity and X velocity. In the projections to X velocity and Z coordinates as well as Y angular velocity and X velocity, the minimum attention slightly regularizes the localized peak. In the projection to Z velocity and X velocity, the minimum attention makes the feedback reinforced and spread out. Heatmaps of policy sensitivity ($\partial u/\partial x||^{2}$, log scale for (a)) for HalfCheetah, comparing vanilla MB-MPO and our method with minimum attention. (a) Torso height vs. forward velocity: Our method learns a simpler strategy with a stable base posture, reducing sensitivity at high speeds. (b) Forward velocity vs. vertical velocity: Our method distributes sensitivity over a wider range of vertical velocities, avoiding sharp reactive hotspots. (c) Forward velocity vs. angular velocity: Our method induces a proactive stabilization mechanism, showing higher sensitivity to torso roll at low speeds.
@@
  • In locomotion (Figure 6a, height vs. velocity): The vanilla policy learns a complex, reactive strategy with high sensitivity at high speeds. Our regularization encourages a simpler, more robust base posture, drastically reducing sensitivity and demonstrating a more efficient control strategy.
  • In vertical motion (Figure 6b, velocity vs. Z-velocity): The vanilla policy relies on a concentrated, high-magnitude ”hotspot” of sensitivity to manage the takeoff phase. Our method distributes this control effort over a wider range of vertical velocities, replacing the sharp, reactive correction with a smoother, more continuous control function.
  • In stabilization (Figure 6c, velocity vs. angular velocity): A key strategic difference is revealed. While the vanilla policy is largely passive to torso roll at low speeds, our regularized policy learns a proactive stabilization mechanism, actively applying corrections in this vulnerable regime to prevent future instability.
Figure 7: Heatmaps of policy jerk ($\partial u/\partial t||^{2}$) for HalfCheetah. Our method concentrates propulsive jerk and introduces a novel, proactive stabilization behaviour against torso roll at low speeds.
Figure 7 shows a similar analysis for temporal sensitivity (jerk). The regularized policy is seen to concentrate its propulsive jerk into more precise phases of the gait cycle, while again introducing proactive, stabilizing jerk to counter torso roll at low speeds—a behavior absent in the baseline. Taken together, these visualizations show a clear shift from a reactive, sporadic control strategy to a proactive, structured, and efficient one. Similarly, more detailed analysis is provided for Hopper, Walker2D, and Humanoid in Appendix B.
@@
Figure 8: Evolution of spatial sensitivity (feedback norm, $\partial u/\partial x||^{2}$) for HalfCheetah. Top Row (MB-MPO): Converges to a broadly reactive strategy. Bottom Row (Ours): Converges to a sparse, structured strategy, focusing sensitivity on a critical phase.
Figure 9: Evolution of temporal sensitivity (feedforward norm/jerk, $\partial u/\partial t||^{2}$) for HalfCheetah. Top Row (MB-MPO): Progresses from chaotic flailing to a moderately jerky running gait. Bottom Row (Ours): The regularizer shapes the policy to be smoother overall, concentrating its action changes into a single, precise phase of the gait cycle.
@@
Maturation of learned control strategies
To understand how the final control strategies emerge, we visualize the evolution of both spatial sensitivity (feedback/Jacobian norm) and temporal sensitivity (feedforward/jerk) at different stages of training. Figures 8 and 9 show the policy heatmaps for HalfCheetah at 10K, 100K, and 200K epochs for both vanilla MB-MPO and our regularized method.
Maturation of feedback (Jacobian) strategy (Figure 8): At the early stage of training (10K epochs), both policies exhibit broad, unstructured sensitivity. As training progresses, a clear divergence in strategy appears. The vanilla MB-MPO policy (top row) develops a wide, high-sensitivity band at high forward velocities, indicating it has learned a reactive strategy that requires constant, high-gain corrections to maintain a running gait. In contrast, our regularized policy (bottom row) converges to a much sparser and more structured landscape. It learns to be largely insensitive across most of the state space, focusing its sensitivity into a narrow, precise band corresponding to the apex of its running motion. This shows the minimum attention actively guides the policy towards a simpler, more efficient state-action mapping.
Maturation of feedforward (jerk) strategy (Figure 9): The evolution of temporal sensitivity is even more striking. The vanilla MB-MPO policy (top row) begins with extremely high, undifferentiated jerk at low speeds (10K epochs), essentially an exploratory behavior. As it learns, it suppresses this initial jerk but retains a moderately high level of action change during the low-crouch running phase (200K epochs).
Figure 10: Left Column: Phase-space analysis of control effort for Hopper’s balancing motion. Top Row (MB-MPO): A ”panic” strategy with an extreme feedback hotspot (>> 2000) for backward leans. Bottom Row (Ours): A smooth, proactive strategy with distributed, low-gain feedback, eliminating the need for high-gain emergency corrections. Right Column: Control effort during a Hopper trajectory. Top (MB-MPO): Shows massive, periodic feedback spikes at the bottom of each hop. Bottom (Ours): Exhibits a more regular gait with dramatically lower and smoother control effort.
Dynamic consequences in Hopper
These fundamental changes in the control landscape lead to qualitatively different and more stable behaviors. We illustrate this using phase-space and time-series analyses of the Hopper agent in Figure 10. This figure provides both a phase-space portrait of the balancing strategy and a time-series view of the control effort throughout the hopping gait.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
P. Lee gives thanks to Frank C. Park introduding the formalism of minimum attention by Roger Brockett. S. Gupta and P. Lee were partially supported by NIH Data Science and NIMHD grant, 3U54MD013376-04S3. P. Lee was partially supported by Center for Equitable AI and Machine Learning Systems (CEAMLS) Affilated Project (ID: 10072401).
@@ Appendix A Technical Appendices and Supplementary Material
Appendix B Detailed Analysis of Policy Sensitivity Heatmaps
This section provides a detailed interpretation of the policy sensitivity heatmaps for each environment, complementing the summary in the main text. In all figures, the left column corresponds to the vanilla MB-MPO baseline, and the right column corresponds to our method, MB-MPO with Minimum Attention.
B.1 Hopper (Figure S1)
B.2 Walker2D (Figure S4)
B.3 Humanoid (Figure S7)
@@
Average total rewards 5531 $\pm$ 127 6564 $\pm$ 57 4242 $\pm$ 68 5870 $\pm$ 117
@@
Table S3: Hopper: Meta-training
@@
Table S5: Walker2D: Meta-training
@@
Table S7: Humanoid: Meta-training

### v3

@@ 4.2 The details of the algorithm implementation
Figure 2: (a) Training of Half-Cheetah. There is a reduction of variance from minimum attention, (b) Comparison of learning curves among model-free RL (SAC, (Haarnoja et al., 2018)) and model-based RL (MB-MPO, (Clavera et al., 2018), the proposed MB-MPO + minimum attention, weighting factor $\alpha=0.05$ and 11). The experiment is run with 10 episodes.
4.3 Dependency on the ensemble model size and imaginary sample size
Figure S1 shows the influence of the ensemble model size of $M$ and the imaginary sample size of $N$. When the ensemble model size of $M$ increases to 20, the total reward also increases. The imaginary sample size $N$ provides the benefit of increased total reward to $N=128$, but decreases when $N$ is increased to 256.
4.4 Comparison with state-of-the-art RL algorithms
The comparision is mostly based on the performance on standard locomotion tasks. We consider a model-free RL, SAC (Soft Actor-Critic) (Haarnoja et al., 2018) and a model-based RL, MB-MPO (Model-Based Meta-Policy Optimization) (Clavera et al., 2018). Firstly, we show the training of Half-Cheetah (Figure 2) in comparison of learning curves among the model-free RL (SAC), the model-based RL (MB-MPO), and the proposed MB-MPO with minimum attention, where the weighting factor is $\alpha=0.01$, $0.05$, and $1$. We have experienced a statistically significant reduction of variance from minimum attention. To demonstrate the generality of our approach, Figure 3 presents the learning curves for Hopper, Walker2D, and Humanoid. Across all three kinematically distinct agents, our method (MB-MPO + minimum attention, blue) consistently converges to a higher average return than the vanilla MB-MPO baseline (red). For the complex Humanoid task, our method also exhibits noticeably lower variance, indicating a more stable and reliable learning process.
The outperformance of MB-MPO with minimum attention is quantified in numbers in Table 1 and S1. Table 1 provides a detailed quantitative summary of the final performance on HalfCheetah after 200K iterations. In the standard meta-training setting, the best minimum attention factor ($\alpha=1.0$) achieves a 45% higher reward than MB-MPO, while simultaneously reducing the average feedback norm (Jacobian) by 28% and energy consumption by 15%. This demonstrates that the regularization does not trade performance for smoothness, but rather enables the discovery of policies that are both higher-performing and more efficient. Crucially, in the meta-testing scenario where the agent must adapt to a crippled leg, the performance benefits of our method are maintained or even amplified. The half-cheetah with minimum attention achieves a higher reward ($6822\pm 109$) than the vanilla baseline ($6356\pm 132$) and does so with a lower feedback norm and feedforward norm (jerk). This suggests that the smoother policies learned through minimum attention provide a better prior for rapid adaptation to significant out-of-distribution embodiment changes.
@@ 4.4 Comparison with state-of-the-art RL algorithms
We first establish the performance of our method on standard locomotion tasks.
Our approach not only matches the sample efficiency of MB-MPO but also converges to a significantly higher average reward. Notably, the learning curves of our method also exhibit reduced variance across episodes, suggesting a more stable and reliable learning process, a key benefit for safe RL.
@@ 4.5 Analysis of learned control strategies of minimum attention
$\bullet$ In locomotion (Figure 5a, height vs. velocity): The vanilla policy learns a complex, reactive strategy with high sensitivity at high speeds. Our regularization encourages a simpler, more robust base posture, drastically reducing sensitivity and demonstrating a more efficient control strategy.
$\bullet$ In vertical motion (Figure 5b, velocity vs. Z-velocity): The vanilla policy relies on a concentrated, high-magnitude ”hotspot” of sensitivity to manage the takeoff phase. Our method distributes this control effort over a wider range of vertical velocities, replacing the sharp, reactive correction with a smoother, more continuous control function.
$\bullet$ In stabilization (Figure 5c, velocity vs. angular velocity): A key strategic difference is revealed. While the vanilla policy is largely passive to torso roll at low speeds, our regularized policy learns a proactive stabilization mechanism, actively applying corrections in this vulnerable regime to prevent future instability. Figure 6 shows a similar analysis for temporal sensitivity (jerk). The regularized policy is seen to concentrate its propulsive jerk into more precise phases of the gait cycle, while again introducing proactive, stabilizing jerk to counter torso roll at low speeds—a behavior absent in the baseline. Taken together, these visualizations show a clear shift from a reactive, sporadic control strategy to a proactive, structured, and efficient one. Similarly, more detailed analysis is provided for Hopper, Walker2D, and Humanoid in Appendix B.
Figure 5: Heatmap of feedback. With Half-Cheetah, the feedback ($\partial u/\partial x||^{2}$) in the projection of the states to a) X velocity and Z coordinates, b) Z velocity and X velocity, and c) Y angular velocity and X velocity. In the projections to X velocity and Z coordinates as well as Y angular velocity and X velocity, the minimum attention slightly regularizes the localized peak. In the projection to Z velocity and X velocity, the minimum attention makes the feedback reinforced and spread out.
Figure 6: Heatmaps of feedforward ($\partial u/\partial t||^{2}$) for HalfCheetah. Our method concentrates propulsive jerk and introduces a novel, proactive stabilization behaviour against torso roll at low speeds.
@@ 5 Conclusion
Acknowledgements
PL gives thanks to Frank C. Park introduding the formalism of minimum attention by Roger Brockett. SG and PL were partially supported by NIH Data Science and NIMHD grant, 3U54MD013376-04S3. PL was partially supported by Center for Equitable AI and Machine Learning Systems (CEAMLS) Affilated Project (ID: 10072401).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Technical Appendices and Supplementary Material
Table S1: Comparison of the model-based RL (MB-MPO) with regularization of minimum attention in meta-training and testing. Total reward at the epoch of 200K. For Hopper, the thigh is paralyzed. For Walker2D, the right thigh is paralyzed.
Half-Cheetah Hopper Walker2D Humanoid
MB-MPO training 6692 $\pm$ 318 2475 $\pm$ 3.81 2399 $\pm$ 694 575 $\pm$ 177
MB-MPO meta-testing 6356 $\pm$ 132 467 $\pm$ 100 523 $\pm$ 174 315 $\pm$ 67
Figure S1: Half-Cheetah: Dependency on the ensemble model size and imaginary rollouts. When the ensemble model size of $M$ increases to 20, the total reward also increases. The imaginary sample size $N$ provides the benefit of increased total reward to $N=128$, but decreases when $N$ is turned to 256.
Appendix B Detailed Analysis of Meta-RL with Minimum Attention
This section provides a detailed decomposition of minimum attention and interpretation of the policy sensitivity heatmaps for each environment, complementing the summary in the main text.
B.1 Half Cheetah (Figure S1, S2, S3, Table S2, S3)
@@
Average total rewards 5531 $\pm$ 127 6564 $\pm$ 57 $42$ $\pm$ 68 5870 $\pm$ 117
@@
Maturation of learned control strategies
To understand how the final control strategies emerge, we visualize the evolution of both spatial sensitivity (feedback/Jacobian norm) and temporal sensitivity (feedforward/jerk) at different stages of training. Figures S2 and S3 show the policy heatmaps for HalfCheetah at 10K, 100K, and 200K epochs for both vanilla MB-MPO and our regularized method.
  • Maturation of feedback (Jacobian) strategy (Figure S2): At the early stage of training (10K epochs), both policies exhibit broad, unstructured sensitivity. As training progresses, a clear divergence in strategy appears. The vanilla MB-MPO policy (top row) develops a wide, high-sensitivity band at high forward velocities, indicating it has learned a reactive strategy that requires constant, high-gain corrections to maintain a running gait. In contrast, our regularized policy (bottom row) converges to a much sparser and more structured landscape. It learns to be largely insensitive across most of the state space, focusing its sensitivity into a narrow, precise band corresponding to the apex of its running motion. This shows the minimum attention actively guides the policy towards a simpler, more efficient state-action mapping.
  • Maturation of feedforward (jerk) strategy (Figure S3): The evolution of temporal sensitivity is even more striking. The vanilla MB-MPO policy (top row) begins with extremely high, undifferentiated jerk at low speeds (10K epochs), essentially an exploratory behavior. As it learns, it suppresses this initial jerk but retains a moderately high level of action change during the low-crouch running phase (200K epochs).
Figure S2: Half Cheetah: Evolution of spatial sensitivity (feedback norm, $\partial u/\partial x||^{2}$). Top Row (MB-MPO): Converges to a broadly reactive strategy. Bottom Row (Ours): Converges to a sparse, structured strategy, focusing sensitivity on a critical phase.
Figure S3: Half Cheetah: Evolution of temporal sensitivity (feedforward norm/jerk, $\partial u/\partial t||^{2}$). Top Row (MB-MPO): Progresses from chaotic flailing to a moderately jerky running gait. Bottom Row (Ours): The regularizer shapes the policy to be smoother overall, concentrating its action changes into a single, precise phase of the gait cycle.
B.2 Hopper (Figure S4, S5, S6, S7, Table S4, S5)
Table S4: Hopper: Meta-training
@@ B.2 Hopper (Figure S4, S5, S6, S7, Table S4, S5)
Table S5: Hopper: Meta-testing
@@
Dynamic consequences in Hopper
These fundamental changes in the control landscape lead to qualitatively different and more stable behaviors. We illustrate this using phase-space and time-series analyses of the Hopper agent in Figure S7. This figure provides both a phase-space portrait of the balancing strategy and a time-series view of the control effort throughout the hopping gait.
Figure S7: Hopper. Left Column: Phase-space analysis of control effort for Hopper’s balancing motion. Top Row (MB-MPO): A ”panic” strategy with an extreme feedback hotspot (>> 2000) for backward leans. Bottom Row (Ours): A smooth, proactive strategy with distributed, low-gain feedback, eliminating the need for high-gain emergency corrections. Right Column: Control effort during a Hopper trajectory. Top (MB-MPO): Shows massive, periodic feedback spikes at the bottom of each hop. Bottom (Ours): Exhibits a more regular gait with dramatically lower and smoother control effort.
B.3 Walker2D (Figure S8, S9, S10, Table S6, S7)
Table S6: Walker2D: Meta-training
@@ B.3 Walker2D (Figure S8, S9, S10, Table S6, S7)
Table S7: Meta-testing of Walker2D
@@
B.4 Humanoid (Figure S11, S12, S13, Table S8, S9)
Table S8: Humanoid: Meta-training
@@ B.4 Humanoid (Figure S11, S12, S13, Table S8, S9)
Table S9: Meta-testing of Humanoid